### v5

@@ 6 The What and Why of Consciousness
Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem. We started with the observation that a human body self organises to maintain itself in the face of change both within and without. An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment. One can regard them as policies according to which the environment is interpreted. To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
We will now argue that: i) there is something it is like to be an organism with a 1ST order self, ii) that subjective experience is the process of learning and enacting a hierarchy or causal identities, and iii) that access consciousness requires both 1ST and 2ND order selves.
Because we do not presuppose representations, we explain how they emerge from attraction to and repulsion from contentless physical states. This means the foundation of information processing is valence.
As we saw earlier, learned policy has valence, and it is a classifier of inputs. Those inputs are “information” in the mechanistic sense of the environment being in one state and not another. Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
@@
For biological, living self-organising systems, being alive is intrinsically good, being dead or ill is bad. For example, in developing self-organising systems such as the human body, it is good to have a better developed sense of smell than vision at early stages of life (babies are more accurate smellers than adults). Different senses have different degrees of valences at different stages of the developmental ladder. Hence there is “something what it is like” to be a baby which is different from something what it is like to be an adult. However, the idea here is that basic experiences (i.e. the self-organising process of staying alive) have valence at their very core (good=stay alive; bad=dead). This culminates in experiential, phenomenal ‘quality’ of appearance, of what it is likeness and how thing appear to the system (i.e. phenomena). That is the aspect that philosophers typically tackle. Our aim is not to dismiss the latter; our claim is that one cannot have the latter without the former.
Clearly, prelinguistic self-organising systems must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions. To learn is to sense “something” and, motivated by valence, construct a policy classifying that “something”. When we say physical states are contentless, we are not suggesting physical objects and properties don’t exist. Objectively, a physical object is an aspect of the environment and every aspect is realized by one state or another. However, previous work has shown that for every aspect of the environment there is an equivalent program, meaning from an objective point of view what is or is not a physical object or property is a matter of interpretation [41]. Subjectively an object, property or quality only exists if an organism constructs a causal identity for it. At the most basic level, the foundation of all of this is physical attraction and repulsion. An organism hard-wired by natural selection can be physically attracted or repelled without any concept of what it is physically attracted or repelled from. It doesn’t know to what it is attracted or repelled from in the sense that it has not constructed a causal identity for that object. There is no named object which has the property ‘attractive’. The organism itself is simply attracted or repelled according to hard-wired policy. For example, a single-celled organism might be attracted to glucose, and its response to ‘tumble’ or ‘stop’ would be hard-wired. It does not satisfy the incentive or scale preconditions to construct a causal identity for glucose, or ‘sweet’, or hunger, or anything else. Intuitively, one might think of this like ‘one dimensional’ valence. To the bacteria, there is no subjective categorical variable or object to which it is attracted, though in the objective sense such a thing does exist. It is simply attracted or repelled. Over the course of this paper we have sought to explain how we get from simple ‘one dimensional’ hard-wired valence to a property or quality that has valence. Categorical variables for properties or ‘qualities’ like ‘the colour red’ or ‘the smell of coffee’ only exist when we have a causal identity for them, which requires scale and incentive. It can be hard-wired, or learned. When we have only one cell we do not have a vocabulary capable of satisfying the scale precondition. However self-organising biological systems are collective [78]. Control is delegated and distributed among smaller components [43] which operate concurrently. An individual cell might only have ‘one dimensional’ valence, but as soon as we have two cells we have something analogous to a second ‘dimension’. Cellular collectives are polycomputational systems, meaning one cell can play a part in more than one computation simultanously, at different scales [79] [^15] This is what the weakness of a policy captures; the number of interactions of which it can be a part.. As we scale up the collective, its is impelled by the cells of which it is formed. Instead of ‘one dimensional’ valence, we now have a rich tapestry of competing drives. It is this rich tapestry that is a property or ‘quality’ of a state of the environment interpreted by the organism. It is this rich tapestry that can be reduced to causal identities for properties like ‘the color red’, ‘hunger’ or ‘thirst’. A higher level of abstraction is formed from this rich tapestry of valence and how it changes over many interactions (the different inputs and outputs in a task that imply generalizable causal identities). Hunger and thirst might have the same overall intensity and so one could say they have the same overall valence, but they are qualitatively different because they are different ‘tapestries’ of valence at a lower level of abstraction. This is what is meant by The Psychophysical Principle of Causality in section 3. In a self-organising system of sufficient scale with sufficient incentive, what starts out as simple ‘one dimensional’ valence culminates in categorical variables that serve different organismic needs. This is why the sound of middle C is different from the color blue. They have different causal identities. Middle C results from sound waves hitting the ear, blue results from light waves hitting the eye. These variations produce different qualia. Qualia do not need to be perfect. They simply need to be “good enough” to motivate actions necessary for survival in the environment of the organism. Qualia are produced by the sensory and nervous systems of the body. The valence of red light waves misses the eyes of dogs entirely. Dogs rely more on smell than sight. This suits their environmental niche. At a sufficient scale causal identities can be constructed, stored and recalled by their quality and valence.
@@
Experience in the sense of experiential phenomenological content cannot exist without experience in the sense of embodied exploration of the body and world as we defined here (see also Ciaunica in prep). To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality. Quality comes first and experiences should be regarded on a continuum rather than a switch on/ switch off phenomena. All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep). One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being. One is already consciously experiencing the world before one can explicitly access one’s own experiences in 2ND order selves.
@@
It follows that every policy learned in this way must classify a quality. Hence every such policy is a local state. A causal identity imbued with disposition by valence. There would be a policy for one’s act of smelling coffee. For perceiving one’s friend. There would be something different it is like to interact with a hostile version of that very same friend. The 1ST order self accompanies everything an organism does. It has a quality, so the 1ST order self is “what it is like” to be that organism. Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [1]. An organism does not make a decision to interpret information, it just reacts. What we call subjective experience is the process of learning and enacting a lattice of causal identities, and there is something that “has” all of those subjective experiences once there is a 1ST order self that is part of them all. This is where phenomenal consciousness begins, with experience as learning and exploration.
@@
One’s 2ND order selves would also have a certain qualitative character, and one’s 3RD order too.
@@
There would, however, be a very clear delineation between conscious and not. The absence of a 1ST order self would mean there would be no policy linking all interventions together, and so no “self” that experiences making them. Hence a 1ST order self must precede higher orders.
@@
Our arguments align with those of Merker, who has linked subjective experience to reafference [80, 71, 72]. We agree that reafference is key, but we provide a very different explanation of why and how. Their work presents biological evidence for subjective experience in organisms with reafference. In contrast, we derive the 1ST order self from first principles and explain why and how it is a classifies “what it is like” to be a particular organism. The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises. Hence these are complementary positions.
7 Unifying Lower and Higher Order Theories of Consciousness
The psychophysical principle of causality takes us from attraction to and repulsion from physical states, to higher order meta representations. This produces different orders of self. We suggest that the full richness of human subjective experience depends upon these different orders of self [^16] Perhaps the full richness of “hard” consciousness [81] depends upon the interaction of selves, perhaps in planning [82]..
@@ 7 Unifying Lower and Higher Order Theories of Consciousness
Conscious “access” as a human has it involves meaningful report, which as we’ve established requires 2ND order selves. This is very different from the mere “access” to information. However, supposing we could somehow contrive a system with 2ND order selves but no 1ST, then 2ND order selves alone would still not allow one to reason about how one’s interventions might affect the contents of 2ND order selves to communicate meaning. For that, one requires a 1ST order self. Hence access consciousness requires both 1ST and 2ND order selves.
Our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational. In pancomputational enactivism there are only contentless states and the programs they imply. Any abstract “representational” content is just organised phenomenal content, clustered according to what causes valence. We don’t discard these two sorts of consciousness, but unify them by showing how phenomenal consciousness gives rise to access.
Our framework thus makes a philosophical zombie impossible because there can be no perfect unconscious replica of a conscious human. The hard problem has it backwards. The question is not why qualia exist, but why anyone thinks representational contents can exist without first being learned through qualitative experience and discrimination.
Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour. Cells are a material with agency [83], and human intelligence is the high level goal directed behaviour of a swarm of cells [78, 84, 28]. When we embody human abstractions (e.g. arithmetic) in silico (e.g. the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [43]. We call the information embodied in the computer “representational” because it means something to us. However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [41]. This suggests there is no such thing as representational contents. They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature: the causes of valence. This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
@@ 8 From Rocks to Einstein: The Hierarchy of Being
To illustrate how our argument applies in the real world we describe stages of conscious organism. Each stage follows from scaling up supply and demand for w-maxing, through natural selection. For each stage we point out animals which are likely to be at least so conscious:
@@
Stage 1: Hard Coded
Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist [85] in a stable environment.
@@
Stage 2: Learning
Stage two introduces learning. To learn an organism must store, classify and order historical examples by valence. However there is not something it is like to be stage two, because there is no locus of “self”. A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora. Even Tripedalia cystophora was recently shown to be capable of associative learning [86]. An entirely distributed control system can “learn”. Likewise, stage two is exemplified by nematodes [87, 88] with a centralised nervous system and some ability to adapt with experience. However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities [72]. When starved C. elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [72, 89, 90], whereas something like a bee or an ant can recall and navigate to previously discovered food [91, 92, 93].
@@
Stage 3: 1ST Order Self
This is where phenomenal consciousness begins, with a 1ST order self. In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have [80, 71, 72]. They identified a housefly as a good example of where subjective experience may begin, and we concur. We also hold this is where an organism might be said to have intent. Intuitively, the policy that motivated behaviour is the intent of that behaviour [^17] In the same way declarative and imperative programs are equivalent [42].. For example, eating tends to involve the intent of satisfying hunger.
A stage three organism can feel simple things like hunger, but cannot cannot conceive of itself from another’s perspective. Subsequently it cannot communicate in the Gricean sense [75, 56], or conceive of its own death, or experience shame.
  • What: A 1ST order self. Reafference. Phenomenal consciousness.
@@
Stage 4: 2ND Order Selves
Stage four is the 2ND order self, and this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense. The ability of ravens to intentionally deceive [94] suggests they are at least stage four. Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$. $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$. It seems likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived. For example, a cat anticipates its prey will flee when it is observed, and hides.
  • What: Access consciousness. Theory of mind. Self-awareness. Inner narrative.
@@
  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent. This could be anticipating the reactions of predator or prey, or navigating a social hierarchy.
@@
Stage 5: 3RD Order Selves
A 3RD order self facilitates complex planning, particularly of a social nature. A 3RD order self is a 2ND order self for one’s 2ND order self. Humans appear to possess this, because we are aware that we are aware. Altruistic behaviour observed in Australian magpies [95] suggests they may also have 3RD order selves.
  • What: Meta self-awareness. Inner narrative in which actors have inner narratives.
@@
Figure 4: Overview of stages and orders of self.

### v6

@@ 6 The What and Why of Consciousness
Up to now we have developed the conceptual toolbox that we use to cross the “explanatory gap” and explain consciousness from an embodied perspective. We started with the observation that a human body is a system that self-organises to maintain itself in the face of change both within and without its finite boundaries. To achieve this vital goal, the embodied system needs to process information qualitatively, meaning it is intrinsically valanced in service of survival. An organism does not spring into existence “objectively” by understanding objects and properties, but rather constructs them through subject-oriented interaction with the environment. One can regard objects and properties as policies according to which the environment is subjectively interpreted. To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change. But how do we get from these embodied subjective experiences to the subjectivity that is usually linked to phenomenal consciousness in awake adult minds?
We will now argue the following.
  1. 1.
    There is something it is like to be an organism with a 1ST order self.
  2. 2.
    Subjective experience is the process of learning and enacting a hierarchy of causal identities.
  3. 3.
    Access consciousness requires both 1ST and 2ND order selves.
@@
Because we do not presuppose representations, we explain how qualitative interpretations emerge from attraction to and repulsion from contentless physical states. This means the foundation of information processing in biological self-organising systems is valence. For biological, living self-organising systems, being alive is intrinsically good, being dead or ill is bad. For example, in developing self-organising systems such as the human body, it is good to have a better developed sense of smell than vision at early stages of life (babies are more accurate smellers than adults Schaal et al. (2020)). Different senses have different degrees of valences at different stages of the developmental ladder. Hence there is “something what it is like” to be a baby which is different from something what it is like to be an adult. However, the idea here is that basic experiences (i.e. the bodily information processing subserving the ultimate goal of staying alive) have valence at their very core (good=stay alive; bad=dead). This culminates in experiential, phenomenal ‘quality’ of appearance, of what it is likeness and how thing appear to the system (i.e. phenomena). That is the aspect that philosophers typically tackle. Our aim is not to dismiss the latter; our claim is that one cannot have the latter without the former.
As we saw earlier, learned policy has valence, and it is a classifier of inputs. Those inputs are what one may call “information” in the mechanistic sense of the environment being in one state and not another. Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
Clearly, prelinguistic self-organising systems must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions. To learn is to sense “something” and, motivated by valence, construct a policy classifying that “something”. When we say physical states are contentless, we are not suggesting physical objects and properties don’t exist. Objectively, a physical object is an aspect of the environment and every aspect is realized by one state or another. However, previous work has shown that for every aspect of the environment there is an equivalent program, meaning from an objective point of view what is or is not a physical object or property is a matter of interpretation Bennett (2024c). The radical and novel claim here is that, subjectively, an object or property only exists to the extent that that organism constructs a causal identity for it.
At the most basic level, the foundation of all of this is physical attraction and repulsion. An organism hard-wired by natural selection can be physically attracted or repelled without any concept of what it is physically attracted or repelled from. The system doesn’t know to what it is attracted or repelled from in the sense that it has not constructed a causal identity for that object. There is no named object which has the property ‘attractive’. The organism itself is simply attracted or repelled according to hard-wired policy. For example, a single-celled organism might be attracted to glucose, and its response to ‘tumble’ or ‘stop’ would be hard-wired. It does not satisfy the incentive or scale preconditions to construct a causal identity for glucose, or ‘sweet’, or hunger, or anything else. Intuitively, one might think of this like ‘one dimensional’ valence. To the bacteria, there is no subjective categorical variable or object to which it is attracted, though in the objective sense such a thing does exist. It is simply attracted or repelled.
Over the course of this paper we have sought to explain how we get from simple ‘one dimensional’ hard-wired valence to a property or quality that has valence. We suggest that categorical variables for properties or ‘qualities’ like ‘the colour red’ or ‘the smell of coffee’ only exist when we have a causal identity for them, which requires scale and incentive. It can be hard-wired or learned. When we have only one cell we do not have a vocabulary capable of satisfying the scale precondition. However self-organising biological systems are collective Fields and Levin (2020). Control is delegated and distributed among smaller components Bennett (2024a) which operate concurrently. An individual cell might only have ‘one dimensional’ valence [^20] Debatably, as even single cells have complex sub-components that interact with each other., but as soon as we have two cells we have something analogous to a second ‘dimension’. Cellular collectives are polycomputational systems, meaning one cell can play a part in more than one computation simultaneously, at different scales Bongard and Levin (2023).
As we scale up the collective, it is impelled by the cells of which it is formed. Instead of ‘one dimensional’ valence, we now have a rich tapestry of competing drives Bennett (2025a) [^21] The contributing parts of this tapestry are co-instantiated in proper time Bennett (2026a) and causally influence one another in space Bennett (2026c). This is important to understand how it relates to notions of conscious “unity” and “integration” in other theories Dennett and Kinsbourne (1992); Varela et al. (2001); Singer and Gray (1995); Fries (2005); Baars (1988); Dehaene and Naccache (2001); Tononi (2004); Lamme (2006) and psychophysics Eagleman and Sejnowski (2000); Pöppel (1997); Vroomen and Keetels (2010); Wallace and Stevenson (2014).. Our position is that it is this rich tapestry that is a property or ‘quality’ of a state of the environment interpreted by the organism. It is this rich tapestry that can be explained through causal identities for properties like ‘the color red’, ‘hunger’ or ‘thirst’. A higher level of abstraction may be formed from this rich tapestry of valence and how this changes over many interactions (the different inputs and outputs in a task that imply generalizable causal identities). For example, hunger and thirst might have the same overall intensity and so one could say they have the same overall valence, but they are qualitatively different because they are different ‘tapestries’ of valence at a lower level of abstraction. This is what is meant by The Psychophysical Principle of Causality in Section 3.
In a self-organising system of sufficiently complex scale with sufficient incentive [^22] Here we refer to the scale and incentive preconditions formally defined earlier., what starts out as simple ‘one dimensional’ valence culminates in categorical variables that serve different organismic needs. This is why the sound of middle C is phenomenologically different from the color blue. They have different causal identities. Middle C results from sound waves hitting the ear, blue results from light waves hitting the eye. These variations produce different subjective embodied experiences.
@@
To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality. Quality comes first and experiences should be regarded on a continuum rather than a switch on / switch off phenomena. All living systems experience the world through their bodies and as such, there is something it is like to experience the world in that basic way (even when one is asleep). One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being. One is already consciously experiencing the world before one can explicitly access one’s own experiences in 2ND order selves.
@@
It follows that every policy learned in this way must classify a quality. Hence every such policy is a local state. A causal identity is inherently dispositional. It is a product of valence in self-organising systems such as human bodies. There would be a policy for one’s act of smelling coffee. For perceiving one’s friend. There would be something different it is like to interact with a hostile version of that very same friend. The 1ST order self accompanies everything an organism does. It has a quality, so the 1ST order self is “what it is like” to be that organism. Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self Nagel (1974). An organism does not make a decision to interpret information, it just does it.
What we call subjective experience is the process of learning and enacting a lattice of causal identities, and there is something that “has” all of those subjective experiences once there is a 1ST order self that is part of them all. This is where phenomenal consciousness begins, with subjective embodied experience as learning and exploration.
@@
Likewise one’s 2ND order selves would have a certain qualitative character, and one’s 3RD order too.
@@
There would, however, be a very clear delineation between the presence and absence of conscious subjective experience. The absence of a 1ST order self would mean there would be no policy linking all interventions together, and so no “self” that experiences making them. Hence a 1ST order self must precede higher order selves and not the other way around. Our approach thus stays in stark contrast with Higher Order Thought Theories of Consciousness (HOT) Fleming (2019).
@@
Rather our arguments align with those of Merker, who has linked subjective experience to reafference Merker (2005, 2007); Barron and Klein (2016). We agree that reafference is key, but we provide a very different explanation of why and how. Their work presents biological evidence for subjective experience in organisms with reafference. In contrast, we derive the 1ST order self from first principles and explain why and how it is a classifies “what it is like” to be a particular organism. The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises. Hence these are complementary positions.
7 Collapsing The Abstraction Layers
@@
Our theory of embodied consciousness places the subjective experience of embodied organisms in the physical world and not in some parallel, dualist world. This means roughly that the psychophysical principle of causality takes us from attraction to and repulsion from physical states, to higher order meta representations. This produces different orders of self. We suggest that the full richness of human subjective experience depends upon these different orders of self [^23] Perhaps the full richness of “hard” consciousness Boltuc (2012) depends upon the interaction of selves, perhaps in planning Shanahan (2012).. Conscious “access” as a human has it involves meaningful report, which as we’ve established requires 2ND order selves. This is very different from the mere “access” to information, in the sense of impinging signals upon the system or stored in memory. However, supposing we could somehow contrive a system with 2ND order selves but no 1ST, then 2ND order selves alone would still not allow one to reason about how one’s interventions might affect the contents of 2ND order selves to communicate meaning. For that, one requires a 1ST order self. Hence access consciousness requires both 1ST and 2ND order selves.
Our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. Our framework thus makes a philosophical zombie impossible because there can be no perfect unconscious replica of a conscious embodied organism but without the qualitative aspect of information processing attached to it [^24] Removing the qualitative aspects would introduce unnecessary extra steps and thus inefficiencyBennett (2025a).. In a self-organising system is causality all the way down. Layer upon layer of abstraction classifies the causes of valence across spatial and temporal scales, from simple one-dimensional signals up to complex tapestries of valence that are the phenomena that make up a self-organising system’s experience.
Related work argues qualia are nothing more than the tapestries of valence we describe Bennett (2025a). Put another way, it claims qualia are reducible to valence. If one were to subscribe to this view, then one might conclude it dissolves hard problem of consciousness, because the formalism [^25] Stack Theory, and Pancomputational Enactivism within it. describes every conceivable world and implies a philosophical zombie with access but not phenomenal consciousness is impossible in all of them. A philosophical zombie acts exactly as a human but without qualia. If one takes qualia to be tapestries of valence, then a system with a 1ST-order-self and the efficient, qualitative information processing we have described would have qualia. Thus a human without qualia would be less efficient than its conscious counterpart and thus not strictly speaking a philosophical zombie.
@@ 8 From Rocks to Einstein: The Hierarchy of Being
In the last section of the paper we suggest the problem of phenomenal consciousness has it backwards. The question is not why qualia exist, but why anyone thinks representational contents can exist in a prelinguistic organism without direct modelling of qualities (i.e. qualia) and discrimination via information processing of cause and affect. To illustrate how our argument applies in the real world we need to go back to the simplest basic level and describe stages of conscious organism. Each stage follows from scaling up supply and demand for w-maxing, through natural selection. For each stage we point out animals which are likely to be at least so conscious:
@@
Stage 1: Hard coded.
Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist Heylighen (2022) in a stable environment.
@@
Stage 2: Learning.
Stage two introduces learning. To learn an organism must store, classify and order historical examples by valence. However there is not “something it is like” to be stage two, because there is no locus of “self”. A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora. Even Tripedalia cystophora was recently shown to be capable of associative learning Bielecki and others (2023). An entirely distributed control system can “learn”. Likewise, stage two is exemplified by nematodes Yu and Rankin (2022); Willett et al. (2018) with a centralised nervous system and some ability to adapt with experience. However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities Barron and Klein (2016). When starved C. elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” Barron and Klein (2016); Lüersen et al. (2014); Artyukhin et al. (2015), whereas something like a bee or an ant can recall and navigate to previously discovered food Wehner (2013); Seeley (1995); Oades and Isaacson (1978).
@@
Stage 3: 1ST order self.
This is where phenomenal consciousness begins, with a 1ST order self. In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have Merker (2005, 2007); Barron and Klein (2016). They identified a housefly as a good example of where subjective experience may begin, and we concur. We also hold this is where an organism might be said to have intent. Intuitively, the policy that motivated behaviour is the intent of that behaviour [^26] In the same way declarative and imperative programs are equivalent Howard (1980).. For example, eating tends to involve the intent of satisfying hunger, which in turn satisfy the basic policy ‘stay alive’. A stage three organism can feel simple things like hunger, but cannot conceive of itself from another’s perspective. Subsequently it cannot communicate in the Gricean sense Grice (1969); Bennett (2023b), or conceive of its own death, or experience shame.
  • What: 1ST order self. Reafference. Phenomenal consciousness.
@@
Stage 4: 2ND order selves.
Stage four is the 2ND order self, and this is where Block’s reportable access consciousness begins in our formal account, because information is now available for report in the Gricean sense. In the ESM we prove that robust Gricean report requires an audience model, which is exactly a 2ND order self. The ability of ravens to intentionally deceive Bugnyar and Kotrschal (2002) suggests they are at least stage four. Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$. $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$. It seems likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived. For example, a cat anticipates its prey will flee when it is observed, and hides.
  • What: Reportable access consciousness. Theory of mind. Self-awareness.
@@
  • Why: A 2ND order self is necessary to infer, communicate and manipulate intent.
@@
Stage 5: 3RD order selves (narrative selves).
A 3RD order self is a narrative self. It lets an organism treat itself as persisting across time, with a story that links past interventions to future commitments. Formally, a 3RD order self is a 2ND order self for one’s 2ND order self. In the ESM we give a minimal trust and commitment game that separates 2ND from 3RD order selves. It shows that even perfect audience modelling cannot sustain trust unless a publicly legible self binding move exists. Humans appear to possess this. Altruistic behaviour observed in Australian magpies Crampton et al. (2022) suggests they may also have narrative selves.
  • What: Impelling narrative self. Meta self-awareness. Trust.
@@
Hierarchy of Selves 0TH. No self
No stable self tag. Behaviour is reactive and not model based.
1ST. Self tag
Separates self generated interventions from matched observations. This is reafference. Grounds phenomenal consciousness.
2ND. Audience model
Models how an audience will decode the agent. Enables Gricean report. Grounds reportable access consciousness.
3RD. Narrative self
Stabilises an identity across time. Enables long horizon planning, self binding, trust, and an impelling narrative.
NTH. Order selves
Higher order selves may facilitate as yet unexplored capabilities.
Figure 1: Overview of orders of self as nested causal identities. 0TH is an embodied agent without an explicit self tag. 1ST adds a reafference like separator between self generated interventions and matched observations. On our account this grounds minimal phenomenal consciousness. 2ND adds a model of how an audience will decode the agent. This enables Gricean report and grounds reportable access consciousness. 3RD is a narrative self that stabilises an identity across time. This enables long horizon planning, self binding, trust, and an impelling narrative. Formal definitions and proofs are in the ESM, where we derive these orders from generalisation optimal learning under the scale and incentive preconditions. Picture kindly provided by Ricard Solé of The Santa Fe Institute.